I just don't understand this.
In the following system of linear equations, state how to combine the two equations to obtain one equation with only one variable.
2x - 7y = 6
x + 3y = 2

Answer by checkley71(8403)

X+3Y=2 OR X=2-3Y NOW REPLACE X IN THE OTHER EQUATION WITH(2-3Y) & SOLVE FOR Y.
2(2-3Y)-7Y=6
4-6Y-7Y=6
-13Y=6-2
-13Y=4
Y=4/-13
Y=-4/13 ANSWER.
X+3*-4/13=2
X-12/13=2
X=12/13+2
X=2 12/13 OR 38/13 ANSWER.